The sound intensity level of a dog’s bark is 40 dB. The intensity of a rock concert is 10,00 times that of the dog’s bark. What is the sound intensity level of the rock concert
Answer #1

40 dB $- 2 I = 10 loge Do 40- In ly ( 17 ) (4) = log ( 2 20 La w/m² Pirock Gresst = 10²x(108) = los w/m2 los - ßr lolog 70 dB